# From random walks to epidemic spreading: Compartment model with mortality for vector transmitted diseases [^1] PROCEEDINGS ICDEA 2024

### v1

@@ Abstract
Epidemic compartment models have become a popular fashion for describing the propagation of an epidemic. To combine compartment models with random walk simulations gives valuable insights into the microscopic mechanisms of the spreading. Here we focus on the propagation of vector-transmitted diseases in complex networks such as Barabási-Albert (BA) and Watts-Strogatz (WS) types. The class of such diseases includes Malaria, Dengue (vectors are mosquitos), Pestilence (vectors are fleas), and many others. There is no direct transmission of the disease among individuals. Individuals are mimicked by independent random walkers and the vectors by the nodes of the network. The walkers and nodes can be either susceptible (S) or infected and infectious (I) representing their states of health. Walkers in compartment I may die from the infection (entering the dead compartment D) whereas infected nodes never die. This assumption is based on the observation that vectors do not fall ill from their infection. A susceptible walker can be infected with a certain probability by visiting an infected node, and a susceptible node by visits of infected walkers. The time spans of infection of walkers and nodes as well as the survival time span of infected walkers are assumed to be independent random variables following specific probability density functions (PDFs). We implement this approach into a multiple random walkers model. We establish macroscopic stochastic evolution equations for the mean-field compartmental population fractions and compare this dynamics with the outcome of the random walk simulations. We obtain explicit expressions for the basic reproduction numbers $R_{M},R_{0}$ with and without mortality, respectively, and prove that $R_{M}<R_{0}$. For $R_{M},R_{0}>1$ the healthy state is unstable, and the disease is starting to spread in presence of at least one infected walker or node. For zero mortality, we obtain in explicit form the stable endemic equilibrium which exists for $R_{0}>1$ and which is independent of the initial conditions. The random walk simulations agree well with the mean-field solutions for strongly connected (small world) graph topologies, whereas for weakly connected graph architectures (large world) and for diseases with high mortality the agreement is less well or occurs only after a long observation time. We also investigate the effect of confinement measures on the spreading of the disease. Our model has a wide range of interdisciplinary applications beyond epidemic dynamics, for instance in the kinetics of certain chemical reactions, the propagation of contaminants, wood fires, and population dynamics.
@@ 1 Introduction
Since the breakout of the COVID-19 pandemics, epidemic models have attracted considerable attention. A popular way to tackle epidemic spreading are so-called compartmental models, where the individuals of a population are divided according to their states of health. The first compartmental model traces back to the seminal work of Kermack and McKendrick [1] where an individual is in one of the states (compartments) susceptible (to infection) - S, infected and infectious - I, recovered (immune) - R. Standard models of the SIR class are able to describe several features of certain infectious diseases, which include mumps, measles, rubella and others. However, standard SIR models do not exhibit persistent oscillatory behaviors or spontaneous outbursts, features which are often observed in the evolution of epidemics. One of the first works able to capture periodic behavior goes back to the model of Soper [2].
@@
A further class are stochastic compartmental models combined with random walk approaches [13, 14, 18, 20, 22, 23, 24, 27, 29, 30] and non-exponentially distributed compartmental sojourn times leading to non-Markovian models were considered recently [11, 25, 26, 27].
In the present paper, our aim is to study the spreading of the class of vector transmitted diseases in a large population of individuals (random walkers) moving on complex graphs similar to human mobility patterns in complex environments such as cities, street-, and transportation networks. The walkers are mimicking the individuals of the population, and the nodes represent the vectors. We also account for the mortality of infected individuals (random walkers) and assume absence of mortality for the infected nodes (vectors). The latter assumption is based on the observation that the vectors (for instance mosquitos) do not fall ill during their infection, and therefore may not die from it.
The paper has the following structure. In Sect. 2 we establish a mean field model and obtain macroscopic evolution equations in stochastic and explicit forms. We derive the basic reproduction number $R_{M}$ in presence of mortality and show that $R_{M}\leq R_{0}$, i.e. the basic reproduction number $R_{0}$ without mortality is un upper bound. We show by linear stability analysis that the healthy state is unstable for $R_{R},R_{M}>1$ and stable for $R_{R},R_{M}<1$. For zero mortality ($R_{M}=R_{0}$) a globally stable endemic equilibrium exists for $R_{0}>1$ independent of the initial conditions. In Sect. 3 we compare the outcomes of the mean field model with random walk simulations, where we focus on WS and BA graphs. At the end of the paper, we investigate for zero mortality the effect of confinement measures on the spreading of the disease and the basic reproduction number in the small world BA graph and compare it with the case of the large world WS graph. For more detailed demonstrations and discussions, we refer to our recent article [28].
@@ 2.1 Mean field approach
Now we introduce $Z_{S}(t),Z_{I}(t)$ ($N_{S}(t),N_{I}(t)$) indicating the number of walkers (nodes) in compartments S and I, and $Z_{D}(t)$ the non-decreasing number of walkers (in compartment D) died during their infection up to time $t$. We consider the total number of walkers $Z=Z(0)=Z_{I}(t)+Z_{S}(t)+Z_{D}(t)$ and of nodes $N=N_{I}(t)+N_{S}(t)$ to be constant, independent of time. We account only for mortality of walkers during their infections. No natural birth- and dead processes are considered.
@@
with time-independent constant rate parameters $\beta_{w},\beta_{n}>0$. Eqs. (1) correspond to a predator-prey mechanism where susceptible walkers (nodes) are the prey of the predator, the infected nodes (walkers). ${\cal A}_{w}(t)$ indicates the infection rate of walkers, the susceptible walkers $S_{w}$ may be infected with a certain probability by (visiting) infected nodes $J_{n}(t)$. ${\cal A}_{n}(t)$ stands for the infection rate of nodes, indicating that susceptible nodes $S_{n}(t)$ may be infected with a certain probability by visits of infected walkers $J_{w}(t)$. As mentioned, no direct transmissions among walkers and among nodes can happen. Infections occur with specific transmission probabilities $p_{w,n}$ which determine the transmission rate constants $\beta_{w,n}$. The latter contain additional topological information when considering the spreading on networks.
@@
with their respective PDFs $K_{i}(\tau)=\{K_{I}^{w}(\tau),K_{I}^{n}(\tau),K_{M}(\tau)\}$ and accounting for causality of the infection rates and PDFs with ${\cal A}_{w,n}(t),K_{i}(t)=0$ for $t<0$ taking into account that no infections occur prior to the outbreak of the disease at $t=0$ and for the positivity of $T_{i}=\{t_{I}^{w},t_{I}^{n},t_{M}\}$. An important special case arises for $f(T_{1},T_{2},T_{3})=g_{1}(T_{1})g_{2}(T_{2})g_{3}(T_{3})$ where from the independence of the $T_{i}$ follows the convenient property
@@
In the second equation of this system comes into play the PDF ${\cal K}_{I,M}^{w}(t)=K_{I}^{w}(t)\Phi_{M}(t)+K_{M}(t)\Phi_{I}^{w}(t)$ capturing all exists of walkers from compartment I, in which the term
$b_{r}(t)=\langle\delta(t-t_{I}^{w})\Theta(t_{M}-t_{I}^{w})\rangle=K_{I}^{w}(t)\Phi_{M}(t)$
@@
$b_{d}(t)=\langle\delta(t-t_{M})\Theta(t_{I}^{w}-t_{M})\rangle=K_{M}(t)\Phi_{I}^{w}(t)$
@@
${\cal K}_{I,M}^{w}(t)=b_{r}(t)+b_{d}(t)=-\frac{d}{dt}[\langle\Theta(t_{M}-t)\Theta(t_{I}^{w}-t)\rangle]=-\frac{d}{dt}[\Phi_{I}^{w}(t)\Phi_{M}(t)]$ (8)
@@
$\int_{0}^{\infty}{\cal K}_{I,M}^{w}(t){\rm d}t=-\int_{0}^{\infty}\frac{d}{dt}[\Phi_{I}^{w}(t)\Phi_{M}(t)]{\rm d}t=\Phi_{I}^{w}(0)\Phi_{M}(0)=1$
@@
since the indicator function $\Theta(t_{M}-t)\Theta(t_{I}^{w}-t)=1$ as long the infection persists, i.e. for $t<t_{IM}^{w}=min(t_{M},t_{I}^{w})$ and where we used independence of $t_{I}^{w},t_{M}$. (9) has the interpretation of the persistence probability of the infection given the infection starts at $t=0$. Further, we will need the following quantities
$\begin{array}[]{clr}\displaystyle{\cal R}(t)=\langle\Theta(t-t_{I}^{w})\Theta(t_{M}-t_{I}^{w})\rangle&=\displaystyle\int_{0}^{t}b_{r}(\tau){\rm d}\tau=\int_{0}^{t}K_{I}^{w}(\tau)\Phi_{M}(\tau){\rm d}\tau&\\ \\ \displaystyle{\cal D}(t)=\langle\Theta(t-t_{M})\Theta(t_{I}^{w}-t_{M})\rangle&=\displaystyle\int_{0}^{t}b_{d}(\tau){\rm d}\tau=\int_{0}^{t}K_{M}(\tau)\Phi_{I}^{w}(\tau){\rm d}\tau&\\ \\ \displaystyle{\cal R}(t)+{\cal D}(t)=\int_{0}^{t}{\cal K}_{I,M}^{w}(\tau){\rm d}\tau&=\displaystyle\int_{0}^{t}[b_{d}(\tau)+b_{r}(\tau)]{\rm d}\tau=1-\Phi_{I}^{w}(t)\Phi_{M}(t)&\\ \\ \displaystyle{\cal D}(\infty)+{\cal R}(\infty)=1&\end{array}$ (10)
@@ 2.2 Zero mortality: endemic equilibrium
$\begin{array}[]{clr}\displaystyle\frac{d}{dt}J_{w}(t)&=\displaystyle{\cal A}_{w}(t)-\left\langle\,{\cal A}_{w}(t-t_{I}^{w})\,\right\rangle-J_{w}(0)\left\langle\,\delta(t-t_{I}^{w})\,\right\rangle&\\ \\ \displaystyle\frac{d}{dt}J_{n}(t)&=\displaystyle{\cal A}_{n}(t)-\left\langle\,{\cal A}_{w}(t-t_{I}^{n})\,\right\rangle-J_{n}(0)\left\langle\,\delta(t-t_{I}^{n})\,\right\rangle&\end{array}$ (13)
@@
$f(\infty)=\lim_{\lambda\to 0}\lambda{\hat{f}}(\lambda)$
@@
independent of initial conditions $J_{w,n}(0)$ where we denote the endemic values from now on $J_{w,n}^{e}=J_{w,n}(\infty)$, $S_{w,n}^{e}=S_{w,n}(\infty)=1-J_{w,n}^{e}$. In these relations, we assume that the PDFs $K_{I}^{w,n}(t)$ are such that $\langle t_{I}^{w,n}\rangle<\infty$. One gets straightforwardly
@@
(i) $\beta_{w}\langle t_{I}^{w}\rangle\to\infty$ (while $\beta_{n}\langle t_{I}^{n}\rangle$ finite): $J_{w}^{e}\to 1$ whereas $J_{n}^{e}\to\frac{\beta_{n}\langle t_{I}^{n}\rangle}{1+\beta_{n}\langle t_{I}^{n}\rangle}<1$
(ii) $\beta_{n}\langle t_{I}^{n}\rangle\to\infty$ (while $\beta_{w}\langle t_{I}^{w}\rangle$ finite): $J_{n}^{e}\to 1$ whereas $J_{w}^{e}\to\frac{\beta_{w}\langle t_{I}^{w}\rangle}{1+\beta_{w}\langle t_{I}^{w}\rangle}<1$
@@ 2.3 Spreading condition with mortality
with ${\cal A}_{w}(t)=\beta_{w}c\,e^{\mu t}$ and ${\cal A}_{n}(t)=\beta_{n}b\,e^{\mu t}$ where $a,b,c,$ are ’small’ constants. Note that (18) is such that $S_{w}(t)+J_{w}(t)+d_{w}(t)=1$ and $S_{n}(t)+J_{n}(t)=1$. Plugging this expression for $\mu\geq 0$ into three independent Eqs. of (2), one gets
@@
$G_{M}(\mu)=1-\beta_{n}\beta_{w}\frac{[1-{\bar{K}}_{I}^{n}(\mu)]}{\mu}\frac{[1-{\hat{\mathcal{K}}}^{w}_{I,M}(\mu)]}{\mu}=0$ (20)
where ${\cal K}_{I,M}^{w}(t)=b_{d}(t)+b_{r}(t)=-\frac{d}{dt}[\Phi_{I}^{w}(t)\Phi_{M}(t)]$ is the PDF that governs the end of the infection (either by recovery or by death). In particular
$\frac{[1-{\hat{\cal K}}^{w}_{I,M}(\mu)]}{\mu}=\int_{0}^{\infty}\Phi_{M}(t)\Phi_{I}^{w}(t)e^{-\mu t}{\rm d}t$
is the LT of the persistence probability $\Phi_{M}(t)\Phi_{I}^{w}(t)=\langle\Theta(t_{M}-t)\Theta(t_{I}^{w}-t)\rangle$ of the walker’s infection, i.e. the probability that $t<min(t_{I}^{w},t_{M})$. The healthy state is unstable if there is a solution $\mu>0$ solving (20). In the following the mean sojourn time of a walker in compartment I with mortality
$\langle t^{w}_{IM}\rangle=\langle min(t_{I}^{w},t_{M})\rangle=\frac{[1-{\hat{\mathcal{K}}}^{w}_{I,M}(\mu)]}{\mu}\bigg{|}_{\mu=0}=\int_{0}^{\infty}t{\cal K}_{I,M}^{w}(t){\rm d}t=\int_{0}^{\infty}\Phi_{M}(t)\Phi_{I}^{w}(t){\rm d}t\,\leq\,\int_{0}^{\infty}\Phi_{I}^{w}(t){\rm d}t=\langle t_{I}^{w}\rangle$ (21)
@@
From (21) follows that $\langle t^{w}_{IM}\rangle\leq\langle t_{I}^{w}\rangle$ (equality only for zero mortality) reflecting simply the feature that $t_{IM}^{w}=min(t_{I}^{w},t_{M})\leq t_{I}^{w}$ with again equality for zero mortality $t_{M}=\infty$. On the other hand, we have $G_{M}(\mu)\to 1$ for $\mu\to\infty$, so there is a positive solution $\mu=\mu_{*}$ of $G_{M}(\mu)=0$ only if
@@
Figure 2: We depict $G_{M}(\mu)$ from Eq. (20) for some $R_{M}$ for exponentially distributed $t_{I}^{w,n},t_{M}$ ($\Phi_{I}^{w,n}(t)=e^{-\xi_{I}^{w,n}t}$, $\Phi_{M}(t)=e^{-\xi_{M}t}$) The parameters are $\beta_{w}=0.5$, $\beta_{n}=2$, $\xi_{I}^{w}=1$, $\xi_{I}^{n}=0.5$. The basic reproduction number without mortality ($\xi_{M}=0$) is $R_{0}=2$ with $R_{M}=R_{0}/(1+\xi_{M})$. One can see that (20) has a positive solution for $R_{M}>1$ indicating instability of the healthy state.

### v2

@@ Abstract
We propose a compartmental model for vector-transmitted diseases, such as Malaria and Dengue, spreading over complex networks. Individuals are represented by independent random walkers and vectors by infected nodes. Both walkers and nodes can be susceptible (S) or infected (I). Infected walkers may die (entering the dead compartment D), while infected nodes remain alive. Susceptible walkers can be infected by visiting infected nodes, and susceptible nodes by visits from infected walkers. We derive explicit expressions for the basic reproduction numbers $R_{0}$ (without mortality) and $R_{M}$ (with mortality), proving that $R_{M}<R_{0}$. When $R_{M},R_{0}>1$, the healthy state is unstable, and for zero mortality, an endemic equilibrium emerges. We also study the effects of confinement measures. Simulations align well with mean-field predictions on strongly connected graphs but deviate for weakly connected networks. Our model has various interdisciplinary applications which include the modeling of chemical reaction kinetics, contaminant spread, and wildfire propagation.
@@ 1 Introduction
Since the breakout of the COVID-19 pandemics, epidemic models have attracted considerable attention. A popular way to tackle epidemic spreading are so-called compartmental models, where the individuals of a population are divided according to their states of health. The first compartmental model traces back to the seminal work of Kermack and McKendrick [1], where an individual is in one of the states (compartments) susceptible (to infection) - S, infected and infectious - I, recovered (immune) - R. Standard models of the SIR class are able to describe several features of certain infectious diseases, which include mumps, measles, rubella and others. However, standard SIR models do not exhibit persistent oscillatory behaviors or spontaneous outbursts, features which are often observed in the evolution of epidemics. One of the first works able to capture periodic behavior goes back to the model of Soper [2].
@@
A further class are stochastic compartmental models combined with random walk approaches [11, 12, 13, 14, 15, 16, 17, 18, 19] and non-exponentially distributed compartmental sojourn times leading to non-Markovian models were considered recently [20, 21, 22, 23].
In the present paper, our aim is to study the spreading of the class of vector transmitted diseases in a large population of individuals (random walkers) moving on complex graphs similar to human mobility patterns in complex environments such as cities, street-, and transportation networks. The walkers are mimicking the individuals of the population, and the nodes represent the vectors. Susceptible walkers may be infected by visiting infected nodes, and susceptible nodes by visits from infected walkers. We also account for the mortality of infected individuals (random walkers) and assume absence of mortality for the infected nodes (vectors). The latter assumption is based on the observation that the vectors (for instance mosquitos) do not fall ill during their infection, and therefore may not die from it. A schematic representation of the vector-borne transmission pathway is shown in Fig. 1.
Figure 1: Vector-borne transmission pathway with mortality of infected walkers.
The paper has the following structure. In Sect. 2 we establish a mean field model and obtain macroscopic evolution equations in stochastic and explicit forms. We derive the basic reproduction number $R_{M}$ in presence of mortality and show that $R_{M}\leq R_{0}$, i.e. the basic reproduction number $R_{0}$ without mortality is un upper bound. We show by linear stability analysis that the healthy state is unstable for $R_{R},R_{M}>1$ and stable for $R_{R},R_{M}<1$. For zero mortality ($R_{M}=R_{0}$) a globally stable endemic equilibrium exists for $R_{0}>1$ independent of the initial conditions. In Sect. 3 we compare the outcomes of the mean field model with random walk simulations, where we focus on Watts-Strogatz (WS) and Barabási–Albert (BA) graphs. At the end of the paper, we investigate for zero mortality the effect of confinement measures on the spreading of the disease and the basic reproduction number in the small world BA graph and compare it with the case of the large world WS graph. For more detailed demonstrations and discussions, we refer to our recent article [24].
@@ 2.1 Mean field approach
It has to be pointed out that for some real-world diseases mortality of vectors may be well non-negligible. However, we focus on the class of diseases, where mortality of vectors is not significant as they do not exhibit symptoms of illness during their infection. We also do not account for natural birth and death processes. This means we consider diseases, where the spreading happens on time-scales for which the demographic change of the population numbers can be neglected. There might exist "slowly" spreading real-world diseases, where natural birth and death processes should be taken into account.
We introduce $Z_{S}(t),Z_{I}(t)$ ($N_{S}(t),N_{I}(t)$) indicating the number of walkers (nodes) in compartments S and I, and $Z_{D}(t)$ the non-decreasing number of walkers (in compartment D) died during their infection up to time $t$. We consider the total number of walkers $Z=Z(0)=Z_{I}(t)+Z_{S}(t)+Z_{D}(t)$ and of nodes $N=N_{I}(t)+N_{S}(t)$ to be constant, independent of time. We account only for mortality of walkers during their infections. Natural birth and death processes are not considered in the present study.
@@
A list of the used symbols and quantities is given hereafter and detailed explanations are given in the text.
$R_{0}$ : basic reproduction number without mortality
$R_{M}$ : basic reproduction number with mortality
$S_{w}$ : fraction of susceptible walkers
$J_{w}$ : fraction of infected walkers
$d_{w}$ : fraction of dead walkers
$S_{n}$ : fraction of susceptible nodes
$J_{n}$ : fraction of infected nodes
${\cal A}_{w}$ : infection rate of walkers
${\cal A}_{n}$ : infection rate of nodes
$\beta_{w}$ : infection rate parameter of walkers
$\beta_{n}$ : infection rate parameter of nodes
$t_{I}^{w}$ : random infection time span of walker without mortality
$t_{I}^{n}$ : random infection time span of nodes
$t_{M}$ : random survival time span of infected walker
$t_{IM}^{w}$ : random infection time span of infected mortal walker
$b_{r}$ : defective recovery PDF of walker
$b_{d}$ : defective mortality PDF of walker
$\Phi_{I}^{w}$ : persistence probability of walkers’ infection without mortality
$\Phi_{I}^{n}$ : persistence probability of nodes’ infection
$\Phi_{M}$ : survival probability of infected walker for infinitely long infection time span
${\cal K}_{I,M}^{w}$ : end of infection PDF of mortal walker
${\cal K}_{I}^{w}$ : recovery PDF of walker without mortality
${\cal K}_{I}^{n}$ : recovery PDF of node
${\cal R}$ : cumulative recovery probability of walker
${\cal D}$ : cumulative mortality probability of walker
$\Theta(\ldots)$ : Heaviside unit-step function
$\delta(\ldots)$ : Dirac’s $\delta$-distribution
[…]
${\hat{f}}$ : Laplace transform of $f$
@@
with time-independent constant rate parameters $\beta_{w},\beta_{n}>0$. Eqs. (1) correspond to a predator-prey mechanism, where susceptible walkers (nodes) are the prey of the predator, the infected nodes (walkers). ${\cal A}_{w}(t)$ indicates the infection rate of walkers, the susceptible walkers $S_{w}$ may be infected with a certain probability by (visiting) infected nodes $J_{n}(t)$. ${\cal A}_{n}(t)$ stands for the infection rate of nodes, indicating that susceptible nodes $S_{n}(t)$ may be infected with a certain probability by visits of infected walkers $J_{w}(t)$. No direct transmissions among walkers and among nodes can happen. See again Fig. 1 for a schematic representation of the transmission pathway. Infections occur with specific transmission probabilities $p_{w,n}$ which determine the transmission rate constants $\beta_{w,n}$. The latter contain additional topological information when considering the spreading on networks.
@@
with their respective PDFs $K_{i}(\tau)=\{K_{I}^{w}(\tau),K_{I}^{n}(\tau),K_{M}(\tau)\}$ and accounting for causality of the infection rates and PDFs with ${\cal A}_{w,n}(t),K_{i}(t)=0$ for $t<0$ taking into account that no infections occur prior to the outbreak of the disease at $t=0$ and for the positivity of $T_{i}=\{t_{I}^{w},t_{I}^{n},t_{M}\}$. An important special case arises for $f(T_{1},T_{2},T_{3})=g_{1}(T_{1})g_{2}(T_{2})g_{3}(T_{3})$, where from the independence of the $T_{i}$ follows the convenient property
@@
In the second equation of this system comes into play the PDF ${\cal K}_{I,M}^{w}(t)=K_{I}^{w}(t)\Phi_{M}(t)+K_{M}(t)\Phi_{I}^{w}(t)$ capturing all exits of walkers from compartment I, in which the term
$b_{r}(t)=\langle\delta(t-t_{I}^{w})\Theta(t_{M}-t_{I}^{w})\rangle=K_{I}^{w}(t)\Phi_{M}(t)$ (8)
@@
$b_{d}(t)=\langle\delta(t-t_{M})\Theta(t_{I}^{w}-t_{M})\rangle=K_{M}(t)\Phi_{I}^{w}(t)$ (9)
@@
${\cal K}_{I,M}^{w}(t)=b_{r}(t)+b_{d}(t)=-\frac{d}{dt}[\langle\Theta(t_{M}-t)\Theta(t_{I}^{w}-t)\rangle]=-\frac{d}{dt}[\Phi_{I}^{w}(t)\Phi_{M}(t)],$ (10)
@@
$\int_{0}^{\infty}{\cal K}_{I,M}^{w}(t){\rm d}t=-\int_{0}^{\infty}\frac{d}{dt}[\Phi_{I}^{w}(t)\Phi_{M}(t)]{\rm d}t=\Phi_{I}^{w}(0)\Phi_{M}(0)=1,$ (11)
@@
since the indicator function $\Theta(t_{M}-t)\Theta(t_{I}^{w}-t)=1$ as long the infection persists, i.e. for $t<t_{IM}^{w}={\rm min}(t_{M},t_{I}^{w})$ and we used independence of $t_{I}^{w}$ and $t_{M}$. (12) has the interpretation of the persistence probability of the infection given the infection starts at $t=0$. Further, we will need the following quantities
$\begin{array}[]{clr}\displaystyle{\cal R}(t)=\langle\Theta(t-t_{I}^{w})\Theta(t_{M}-t_{I}^{w})\rangle&=\displaystyle\int_{0}^{t}b_{r}(\tau){\rm d}\tau=\int_{0}^{t}K_{I}^{w}(\tau)\Phi_{M}(\tau){\rm d}\tau&\\ \\ \displaystyle{\cal D}(t)=\langle\Theta(t-t_{M})\Theta(t_{I}^{w}-t_{M})\rangle&=\displaystyle\int_{0}^{t}b_{d}(\tau){\rm d}\tau=\int_{0}^{t}K_{M}(\tau)\Phi_{I}^{w}(\tau){\rm d}\tau&\\ \\ \displaystyle{\cal R}(t)+{\cal D}(t)=\int_{0}^{t}{\cal K}_{I,M}^{w}(\tau){\rm d}\tau&=\displaystyle\int_{0}^{t}[b_{d}(\tau)+b_{r}(\tau)]{\rm d}\tau=1-\Phi_{I}^{w}(t)\Phi_{M}(t)&\\ \\ \displaystyle{\cal D}(\infty)+{\cal R}(\infty)=1,&\end{array}$ (13)
@@ 2.2 Zero mortality: endemic equilibrium
$\begin{array}[]{clr}\displaystyle\frac{d}{dt}J_{w}(t)&=\displaystyle{\cal A}_{w}(t)-\left\langle\,{\cal A}_{w}(t-t_{I}^{w})\,\right\rangle-J_{w}(0)\left\langle\,\delta(t-t_{I}^{w})\,\right\rangle&\\ \\ \displaystyle\frac{d}{dt}J_{n}(t)&=\displaystyle{\cal A}_{n}(t)-\left\langle\,{\cal A}_{w}(t-t_{I}^{n})\,\right\rangle-J_{n}(0)\left\langle\,\delta(t-t_{I}^{n})\,\right\rangle,&\end{array}$ (16)
@@
$f(\infty)=\lim_{\lambda\to 0}\lambda{\hat{f}}(\lambda)$ (19)
@@
independent of initial conditions $J_{w,n}(0)$, where we denote the endemic values from now on $J_{w,n}^{e}=J_{w,n}(\infty)$, $S_{w,n}^{e}=S_{w,n}(\infty)=1-J_{w,n}^{e}$. In these relations, we assume that the PDFs $K_{I}^{w,n}(t)$ are such that $\langle t_{I}^{w,n}\rangle<\infty$. One gets straightforwardly
@@
(i) $\beta_{w}\langle t_{I}^{w}\rangle\to\infty$ (while $\beta_{n}\langle t_{I}^{n}\rangle$ finite): $J_{w}^{e}\to 1$, whereas $J_{n}^{e}\to\frac{\beta_{n}\langle t_{I}^{n}\rangle}{1+\beta_{n}\langle t_{I}^{n}\rangle}<1$
(ii) $\beta_{n}\langle t_{I}^{n}\rangle\to\infty$ (while $\beta_{w}\langle t_{I}^{w}\rangle$ finite): $J_{n}^{e}\to 1$ whereas $J_{w}^{e}\to\frac{\beta_{w}\langle t_{I}^{w}\rangle}{1+\beta_{w}\langle t_{I}^{w}\rangle}<1,$
@@ 2.3 Spreading condition with mortality
with ${\cal A}_{w}(t)=\beta_{w}c\,e^{\mu t}$ and ${\cal A}_{n}(t)=\beta_{n}b\,e^{\mu t}$, where $a,b,c,$ are ’small’ constants. Note that (22) is such that $S_{w}(t)+J_{w}(t)+d_{w}(t)=1$ and $S_{n}(t)+J_{n}(t)=1$. Plugging this expression for $\mu\geq 0$ into three independent Eqs. of (2), one gets
@@
$G_{M}(\mu)=1-\beta_{n}\beta_{w}\frac{[1-{\bar{K}}_{I}^{n}(\mu)]}{\mu}\frac{[1-{\hat{\mathcal{K}}}^{w}_{I,M}(\mu)]}{\mu}=0,$ (24)
where ${\cal K}_{I,M}^{w}(t)=b_{d}(t)+b_{r}(t)=-\frac{d}{dt}[\Phi_{I}^{w}(t)\Phi_{M}(t)]$ is the PDF that governs the end of the infection (either by recovery or by death). In particular,
$\frac{[1-{\hat{\cal K}}^{w}_{I,M}(\mu)]}{\mu}=\int_{0}^{\infty}\Phi_{M}(t)\Phi_{I}^{w}(t)e^{-\mu t}{\rm d}t$ (25)
is the LT of the persistence probability $\Phi_{M}(t)\Phi_{I}^{w}(t)=\langle\Theta(t_{M}-t)\Theta(t_{I}^{w}-t)\rangle$ of the walker’s infection, i.e. the probability that $t<{\rm min}(t_{I}^{w},t_{M})$. The healthy state is unstable if there is a solution $\mu>0$ solving (24). In the following the mean sojourn time of a walker in compartment I with mortality
$\langle t^{w}_{IM}\rangle=\langle{\rm min}(t_{I}^{w},t_{M})\rangle=\frac{[1-{\hat{\mathcal{K}}}^{w}_{I,M}(\mu)]}{\mu}\bigg{|}_{\mu=0}=\int_{0}^{\infty}t{\cal K}_{I,M}^{w}(t){\rm d}t=\int_{0}^{\infty}\Phi_{M}(t)\Phi_{I}^{w}(t){\rm d}t\,\leq\,\int_{0}^{\infty}\Phi_{I}^{w}(t){\rm d}t=\langle t_{I}^{w}\rangle$ (26)
@@
From (26) follows that $\langle t^{w}_{IM}\rangle\leq\langle t_{I}^{w}\rangle$ (equality only for zero mortality) reflecting simply the feature that $t_{IM}^{w}={\rm min}(t_{I}^{w},t_{M})\leq t_{I}^{w}$ with again equality for zero mortality $t_{M}=\infty$. On the other hand, we have $G_{M}(\mu)\to 1$ for $\mu\to\infty$, so there is a positive solution $\mu=\mu_{*}$ of $G_{M}(\mu)=0$ only if
@@
Figure 3: We depict $G_{M}(\mu)$ from Eq. (24) for some $R_{M}$ for exponentially distributed $t_{I}^{w,n},t_{M}$ ($\Phi_{I}^{w,n}(t)=e^{-\xi_{I}^{w,n}t}$, $\Phi_{M}(t)=e^{-\xi_{M}t}$). The parameters are $\beta_{w}=0.5$, $\beta_{n}=2$, $\xi_{I}^{w}=1$, $\xi_{I}^{n}=0.5$. The basic reproduction number without mortality ($\xi_{M}=0$) is $R_{0}=2$ with $R_{M}=R_{0}/(1+\xi_{M})$. One can see that (24) has a positive solution for $R_{M}>1$ indicating instability of the healthy state.